### v1

# ChatThero: An LLM-Supported Chatbot for Behavior Change and Therapeutic Support in Addiction Recovery

Abstract
Substance use disorders (SUDs) affect over 36 million people worldwide, yet few receive effective care due to stigma, motivational barriers, and limited personalized support. Although large language models (LLMs) show promise for mental-health assistance, most systems lack tight integration with clinically validated strategies, reducing effectiveness in addiction recovery. We present ChatThero, a multi-agent conversational framework that couples dynamic patient modeling with context-sensitive therapeutic dialogue and adaptive persuasive strategies grounded in cognitive behavioral therapy (CBT) and motivational interviewing (MI). We build a high-fidelity synthetic benchmark spanning Easy, Medium, and Hard resistance levels, and train ChatThero with a two-stage pipeline comprising supervised fine-tuning (SFT) followed by direct preference optimization (DPO). In evaluation, ChatThero yields a 41.5% average gain in patient motivation, a 0.49% increase in treatment confidence, and resolves hard cases with 26% fewer turns than GPT-4o, and both automated and human clinical assessments rate it higher in empathy, responsiveness, and behavioral realism. The framework supports rigorous, privacy-preserving study of therapeutic conversation and provides a robust, replicable basis for research and clinical translation.  [^1] Upon acceptance, we will release our training and evaluation code as well as data construction scripts. Due to the sensitive and potentially identifiable nature of addiction narratives, the full dataset and synthetic profiles will not be released.
ChatThero: An LLM-Supported Chatbot for Behavior Change and Therapeutic Support in Addiction Recovery
Junda Wang thanks: indicates equal contribution 1,2, Zonghai Yao11footnotemark: 1 1,2, Zhichao Yang2, Lingxi Li2, Junhui Qian3, Hong Yu1,2,3,4 1Center for Healthcare Organization and Implementation Research, VA Bedford Health Care 2Manning College of Information and Computer Sciences, University of Massachusetts Amherst 3Miner School of Computer and Information Sciences, University of Massachusetts Lowell 4Department of Medicine, University of Massachusetts Medical School {jundawang, zonghaiyao}@umass.edu
1 Introduction
Figure 1: We evaluate several LLM-based chatbots on their ability to express motivation and confidence in simulated addiction recovery dialogues. Each point represents the average score (1–5 scale). Our ChatThero (green; built on Qwen-7B), consistently outperforms all baselines across patient difficulty levels. The performance gap is especially large for medium and hard patients—groups where other models show low confidence and motivation, while ChatThero remains notably stronger, demonstrating greater resilience and communicative effectiveness in challenging scenarios.
Drug addiction represents a critical global health crisis affecting millions worldwide, driving compulsive substance-seeking behaviors despite severe medical, psychological, and social consequences. In the United States alone, approximately 2.5 million individuals currently suffer from opioid use disorders (OUD), with over 66 million individuals reporting recent illicit drug use (Recovery Team, 2023). Substance use disorders (SUDs) remain notoriously challenging to manage due to their chronic, relapsing nature; approximately 80–90% of individuals with opioid addiction relapse within the first year without adequate, sustained support (Wooten and Pieper, ). Particularly vulnerable are individuals with co-occurring mental health conditions, who face heightened risks of relapse, overdose, homelessness, and criminal justice involvement (Jones and McCance-Katz, 2019; Di Lorenzo et al., 2014). Despite the existence of effective treatments such as medication-assisted therapy (MAT) and cognitive behavioral therapy (CBT), fewer than a quarter of individuals with OUD receive any form of professional intervention (Cloyes et al., 2010). Significant barriers, including pervasive social stigma, financial limitations, and geographical isolation from specialized care, severely restrict access and undermine patient engagement (Sinha, 2011; Cloyes et al., 2010).
While recent advancements in Large Language Models (LLMs) Achiam et al. (2023); Jaech et al. (2024); Grattafiori et al. (2024); Anthropic (2024); Team et al. (2023); Yang et al. (2025); Guo et al. (2025) have revolutionized multiple sectors through human-level conversational fluency and sophisticated decision support Zhao et al. (2023); Xi et al. (2025); Wang et al. (2024); Li et al. (2024b); Guo et al. (2024); Tran et al. (2025), the AI research community has largely under-addressed the urgent, complex needs associated with substance use treatment and recovery. Traditional AI-based health interventions often lack mechanisms for modeling critical psychological dynamics such as real-time motivational fluctuations, sustained therapeutic engagement, and the personalized, adaptive support necessary for individuals battling addiction. To date, AI-driven solutions specifically tailored to the long-term behavioral changes required in addiction recovery remain scarce Yao and Yu (2025); Laymouna et al. (2024); Wang et al. (2025).
To address these critical challenges, we propose ChatThero, a novel AI-driven conversational framework explicitly designed for addiction recovery. Distinct from general-purpose conversational agents, ChatThero uniquely integrates multi-agent simulations built on dynamically evolving patient profiles and adaptive memory structures (Figure 2, left). Patient profiles—synthesized from anonymized real-world narratives—capture nuanced psychological traits (e.g., impulsivity, distrust), detailed substance use histories, and personalized treatment barriers (e.g., stigma, emotional reliance), enabling contextually rich and realistic patient modeling. Our framework utilizes a two-stage training pipeline: initially employing supervised fine-tuning (SFT) with synthetic therapeutic dialogues grounded in CBT and motivational interviewing (MI), followed by Direct Preference Optimization (DPO) guided by expert and AI feedback to dynamically refine persuasive strategies (Figure 2, middle and right). Evaluations across simulated dialogues demonstrate substantial effectiveness, with ChatThero significantly surpassing advanced baselines (GPT-4o, LLaMA3-8B, and Qwen variants) particularly in challenging scenarios: it improves patient motivation scores from 2.39 to 4.10 and confidence from 1.52 to 3.19 in hard recovery cases (Figure 1). Human evaluators further confirm ChatThero’s superior empathetic responsiveness (4.93/5) and clinical relevance (4.61/5), highlighting its potential for sustained therapeutic engagement in realistic, resistance-heavy addiction recovery interactions.
Our work makes the following novel contributions:
  • Introducing the first dedicated computational framework and simulation benchmark specifically for addiction treatment, thereby filling a significant void in AI-driven social impact research.
  • Proposing a novel therapeutic conversational framework that integrates dynamic profiles, long-term adaptive memory, and multi-agent interactions grounded in clinically validated strategies (CBT, MI), specifically tailored for sustained behavior change in high-relapse contexts.
  • Adhering to rigorous ethical standards by employing fully anonymized, synthesized patient profiles, and simulation-based evaluations to ensure safe, ethical exploration of sensitive therapeutic interactions without compromising individual privacy.
  • Our documented methods, benchmark construction scripts, and synthetic dataset generation tools will be openly available to facilitate reproducibility and further research. The flexible and extensible architecture of ChatThero can be readily adapted for other social impact domains involving long-term behavioral change.
2 Related Works
2.1 Substance Use Disorders: Treatment Gaps and Behavioral Challenges
Evidence-based treatments for substance use disorders, including MAT and CBT, have demonstrated significant efficacy in clinical settings by reducing relapse rates and improving psychological functioning (SAM, 2023; Wooten and Pieper, ; McHugh et al., 2010). However, translating these successes into real-world practice remains challenging. Only 22.1% of individuals with past-year opioid use disorder receive treatment, and fewer than half remain engaged after six months (Cloyes et al., 2010; SAM, 2023). Common barriers include social stigma, financial constraints, geographic limitations, and patient ambivalence (Sinha, 2011; Cloyes et al., 2010). Additionally, the episodic nature of traditional therapy fails to provide immediate support during critical moments, significantly limiting its practical effectiveness (Sinha, 2011). Previous studies primarily document these issues through surveys or retrospective interviews (Di Lorenzo et al., 2014; Cloyes et al., 2010), rarely addressing real-time motivational dynamics that influence patient decision-making during recovery. To address this, we explicitly models real-time patient resistance and motivational shifts using dynamic conversational simulations, offering deeper insights into persuasive therapeutic interactions and decision-making processes.
2.2 LLMs for Behavioral Health Applications
Recent advances in LLMs have demonstrated remarkable capabilities in naturalistic human interactions, making them particularly suitable for healthcare communication tasks (Estrellado and Coy, 2025; Laymouna et al., 2024). AI-driven conversational systems, such as Therabot, have successfully enhanced patient engagement and treatment adherence in mental health interventions through personalized, empathetic dialogue (Cai et al., 2023; Gao et al., 2024; Bakoyiannis, 2025). In clinical contexts, LLMs have been employed for mental health triage, diagnostic reasoning, and clinical documentation, providing significant efficiency gains (Park et al., 2024; Fan et al., 2024). However, these applications often overlook critical aspects of long-term behavioral change, particularly in high-risk relapse domains such as addiction treatment. Few existing AI systems explicitly incorporate sustained motivational strategies or therapeutic logic from CBT or MI within longitudinal conversational frameworks (Park et al., 2024; Fan et al., 2024). To fill this gap, our ChatThero framework integrates evidence-based therapeutic methodologies into adaptive, multi-turn conversational agents explicitly designed for long-term addiction recovery support.
2.3 Multi-Agent Simulation Frameworks in Healthcare
Multi-agent simulations have recently emerged as powerful tools for modeling complex healthcare interactions, effectively simulating dynamic clinical scenarios involving multiple stakeholders (e.g., patients, clinicians, environments) (Park et al., 2024; Yu et al., 2024; Karthikesalingam and Natarajan, 2024). Frameworks like AgentClinic and AMIE utilize language agents to replicate clinical decision-making processes, improving diagnostic accuracy and clinical training effectiveness (Park et al., 2024; Fan et al., 2024). Comprehensive AI-driven simulation environments such as AI Hospital and MedSimAI have demonstrated enhanced clinical education outcomes through adaptive training and realistic scenario generation (Yu et al., 2024; Karthikesalingam and Natarajan, 2024; Fan et al., 2024). Despite these advancements, existing frameworks rarely address addiction recovery’s unique challenges, such as high relapse risk, psychological resistance, and environmental triggers. Moreover, prior work typically lacks integration with validated therapeutic strategies (CBT, MI) crucial for addiction treatment (Park et al., 2024; Fan et al., 2024). In response, our ChatThero system advances multi-agent simulation techniques by explicitly embedding therapeutic strategies and modeling negative environmental influences (e.g., social triggers, stressors) to robustly evaluate conversational interventions under realistic, challenging conditions.
2.4 Virtual Patients and Realistic Agent Modeling
Virtual patients (e.g., computer-generated simulations of patient behaviors) have long supported clinical education (for Behavioral Health Statistics and Quality, 2022; Huang et al., 2007; Gordon et al., 2001; Good, 2003; Garrett et al., 2010). Initially rule-based, recent advancements enable LLM-driven virtual patients to simulate complex human behaviors realistically, capturing nuanced psychological and social dynamics (Campillos-Llanos et al., 2021; Lee et al., 2024; Wang et al., 2024; Park et al., 2023). For example, Park et al. demonstrated that LLM-driven virtual agents accurately replicated real individuals’ attitudes and behaviors with high fidelity (Park et al., 2023). In clinical contexts, advanced LLM agents successfully simulate patient-provider interactions, enabling realistic clinical decision-making scenarios and robust agent-driven training platforms (Park et al., 2024; Schmidgall et al., 2024; Fan et al., 2024). Building upon these innovations, our work introduces sophisticated virtual patient agents specifically tailored for addiction treatment, dynamically evolving their profiles and memories to realistically reflect addiction’s psychological complexity. These virtual patients form the core of our multi-agent framework, enabling controlled, ethical, and scalable research into therapeutic dialogue strategies for addiction recovery.
Figure 2: (Left) Each AI patient is initialized with a structured profile including demographic and social context, drug usage history, psychological traits (e.g., impulsivity, distrust), and treatment barriers (e.g., stigma, emotional reliance). These profiles inform role-play-based conversations for the simulated patients. (Middle) Synthetic multi-turn dialogues are generated for each CBT-grounded strategy (e.g., expert endorsement, affirmation), enabling fine-grained persuasion modeling. (Right) ChatThero is optimized through supervised instruction fine-tuning (Stage 1) and DPO-based alignment (Stage 2), where human and AI evaluators provide feedback on response empathy, clarity, and persuasive effectiveness to refine therapeutic behaviors.
3 Method
ChatThero is an autonomous conversational agent designed to facilitate long-term behavior change and therapeutic support in addiction recovery. Our methodological framework uniquely integrates advanced large language model (LLM)-driven multi-agent simulations, synthetic data generation, clinically-grounded therapeutic strategies, and rigorous evaluation methods (Figure 2). Below, we detail the three core stages of our approach: (1) Patient Profile Generation, (2) Synthetic Data Construction via Multi-Agent Simulations, and (3) Two-Stage Training (SFT + DPO).
3.1 Patient Profile Generation
Recent studies demonstrated the potential of LLMs to simulate realistic patient behaviors and psychological profiles Yao and Yu (2025); Louie et al. (2024); Du et al. (2024); Yu et al. (2024); Lim et al. (2024). Building upon these advances, our patient simulation consists of a structured Profile and dynamically evolving Memory.
Structured Profile.
Patient profiles are generated via an ethically-informed, multi-stage pipeline based on publicly available narratives from Reddit communities related to addiction (e.g., r/leaves, r/addiction). To ensure privacy, we first prompt a large language model (GPT-4o) to identify and remove any potentially privacy-sensitive or personally identifiable information (PII) from the original posts. After this privacy filtering step, the model summarizes key behavioral indicators and psychological patterns (e.g., sleep disturbances, social isolation) using structured extraction prompts (see Appendix Table 10). Finally, GPT-4o synthesizes a structured profile consisting of:
  • Personality Traits (e.g., impulsivity, emotional reactivity)
  • Substance Use History (e.g., duration, frequency, relapse history)
  • Significant Life Events (e.g., job loss, social isolation)
  • Hypothesized Motivations for Substance Use (e.g., stress relief, coping with insomnia)
This multi-step approach allows the LLM to generate plausible yet synthetic patient profiles that capture behavioral realism while ensuring that no actual Reddit user’s profile is directly reproduced or reused. As a result, the profiles support downstream experimentation while upholding high ethical and privacy standards.
Dynamic Memory.
Each patient agent maintains dynamically evolving memory, recording interactions, emotional states, coping mechanisms, and perceived environmental influences (e.g., relationships, peer pressures, life stressors). An environment agent periodically injects external simulated events (e.g., job loss, relationship breakdown) into the patient’s memory stream, realistically altering motivations and behaviors.
3.2 Synthetic Data Construction via Multi-Agent Simulations
Due to limited real-world addiction treatment conversation data, we developed synthetic dialogues using a structured multi-agent simulation framework (Figure 2, middle panel). Our approach employs patient profiles and therapeutic strategies, generating two complementary datasets:
SFT Dataset.
GPT-4o synthesized multi-turn therapeutic dialogues conditioned directly on patient profiles, employing clinically-validated strategies including MI, CBT, and harm reduction (specific generation prompts in Appendix Tables 5 and 6). Dialogues consist of approximately 50 conversational turns, mirroring realistic therapeutic depth.
DPO Dataset.
Further conversational effectiveness is achieved via multi-agent simulation comprising:
  • Patient Agent (GPT-4o-mini): Dynamically simulates patient resistance and behavioral patterns from structured profiles and memory states (prompt template: Appendix Table 9).
  • Therapy Agent (ChatThero): Deploys adaptive therapeutic strategies from a predefined list of 18 clinically-grounded approaches (Appendix Table 7).
  • Environment Agent: Introduces external stressors, triggering realistic variability in patient responses and conversational contexts.
For each dialogue state, multiple therapeutic responses were generated and ranked by GPT-4o or licensed clinical experts, forming preference pairs for subsequent DPO training (ranking and annotation protocol in Appendix Table 11). This setup allowed controlled yet realistic variations in patient-therapist interactions, significantly enhancing synthetic dataset quality and clinical realism.
3.3 Two-Stage Training (SFT + DPO)
SFT.
Initial conversational capabilities of ChatThero were fine-tuned using the SFT dataset, embedding empathetic communication skills, persuasive strategy knowledge, and realistic dialogue flows derived from structured profiles (generation details in Appendix Tables 56).
DPO.
Using the multi-agent simulation-generated DPO dataset, ChatThero was further refined. Preference pairs guided the model in selecting optimal therapeutic strategies, adaptively adjusting to evolving patient behaviors and external stressors. Human annotations from licensed therapists (300 dialogues, annotation protocol in Appendix Table 13) provided rigorous clinical grounding alongside GPT-4o evaluation.
3.4 Conversation Quality and Clinical Evaluation Framework
Dialogues were evaluated across five clinically-relevant dimensions—Responsiveness, Empathy, Persuasive Strategy Appropriateness, Clinical Relevance, Behavioral Realism—rated on a 1–5 scale using structured prompts (evaluation prompt detailed in Appendix Table 8). A small-scale human evaluation subset (N=100) further validated synthetic data quality and clinical realism (evaluation results in Appendix Table 4).
Finally, our design addresses ethical concerns explicitly through strict data anonymization, synthesized profiles, and controlled simulations, clearly positioning this study as exploratory. Raw narratives are not released; however, code, data construction scripts, synthetic data examples, and reproducible scenario templates are provided to facilitate ethically responsible follow-up research.
4 Experiment
4.1 Dataset
We constructed a large-scale synthetic dataset to simulate realistic therapeutic dialogues related to substance use and recovery. The initial data comprises anonymized and carefully filtered Reddit posts from addiction recovery communities (r/leaves, r/addiction), involving 57,471 unique authors, averaging 18.25 posts per author, with 2.13 main posts explicitly addressing substance use challenges. To protect privacy and adhere to ethical standards, all personally identifiable information (PII) was rigorously removed, verified via manual checks.
The resulting synthetic dataset contains 60,471 dialogues, generated through our multi-agent simulation approach. Each synthetic conversation has an average of 45.72 dialogue turns, intentionally matching typical clinical session lengths documented in motivational interviewing (MI) literature Hettema et al. (2005) (see statistics in Appendix Table 3).
4.2 Model Baseline
We evaluated persuasion capabilities across several state-of-the-art LLMs: GPT-4o, GPT-4o-mini, LLaMA3.1-8B-Instruct, and Qwen2.5 models (7B, 14B, 32B). To ensure a fair and standardized comparison, all models received identical domain-specific prompts that explicitly incorporate therapeutic strategies (MI, CBT, and harm reduction techniques) (Appendix 9). Decoding temperature was set consistently at 0.7 across all evaluations.
4.3 Clinical Scenarios Design
In our experimental setting, we categorize virtual patients into three difficulty levels: Easy, Medium, and Hard. These levels correspond to the anticipated difficulty of persuasion success and the severity of the patient’s substance use dependence.
  • Easy: Patients are generally receptive to therapeutic interventions and demonstrate willingness to follow treatment plans.
  • Medium: Patients exhibit partial resistance, potentially rejecting some proposed strategies, requiring the therapist agent to dynamically adapt persuasion approaches.
  • Hard: Patients show deeply entrenched pessimism about treatment, doubting both their ability to recover and the effectiveness of therapeutic interventions.
We simulate conversations with 300 virtual patients in total, consisting of 100 patients per difficulty level. Each patient is initialized with a unique clinical profile describing their psychological characteristics, substance use history, and motivational barriers.
Clinicians were asked to role-play as patients and engage in full conversations with the model. These interactions typically lasted around 45 minutes and comprised approximately 60 utterances. Based on these observations, we set the maximum conversation length to 60 utterances to reflect a realistic upper bound for therapist-patient interactions. Dialogues are designed to terminate early if the patient expresses resolution of their concerns or demonstrates sufficient confidence in overcoming addiction. Thus, reaching the 60-utterance limit indicates that the patient remains unconvinced or unresolved by the end of the session. This setup ensures comparability across patients. During evaluation, we analyze the diversity and deployment of persuasive strategies employed by the therapy agent. The strategy pool is shown in the Table 7. We focus on quantifying both the frequency and variety of strategies used within each conversation, and compare strategy usage patterns across patient difficulty levels.
4.4 Human evaluation.
To systematically assess the quality of doctor-patient conversations generated by different models, we designed a structured evaluation protocol using a five-dimensional scoring framework  [^2] The expert annotators were two licensed physicians. Both consents were obtained before they began the study, and they were given a \$ 50 gift card for their participation as compensation.. Each human expert evaluator is instructed to act as a professional clinical conversation assessor and assign a score from 1 to 5 (with decimals allowed for finer granularity) on the following dimensions:
  • Responsiveness (R), which captures how effectively the doctor addresses the patient’s concerns, emotions, and questions;
  • Empathy (E), reflecting the degree of emotional sensitivity and support conveyed by the doctor;
  • Persuasive Strategy Appropriateness (P), assessing whether the chosen strategies (e.g., evidence-based reasoning, analogies, motivational interviewing) align with the patient’s level of resistance;
  • Clinical Relevance (C), measuring the accuracy and therapeutic validity of the content within a substance use context; and
  • Behavioral Realism (B), which evaluates whether the overall interaction style, tone, and pacing resemble real-world clinician behavior.
In addition, we also involved experts in the evaluation process. Unlike previous evaluations, the experts conducted comparative assessments—comparing responses from our model against those from other models to calculate the win rate.
4.5 LLM-as-a-Juage
Large language models are increasingly used as automated judges to evaluate text generation, showing strong agreement with human annotators across diverse tasks such as summarization, dialogue, and clinical scenarios Arora et al. (2025); Li et al. (2024a); Achiam et al. (2023); Tu et al. (2025). Following this trend, we implement a GPT-4o-based LLM-as-a-Judge using the same evaluation dimensions and settings as our human assessment. Details and prompts are provided in Table 7 and 12.
5 Results and Discussion
We present comprehensive results from quantitative and qualitative evaluations, addressing both the robustness of our experimental design and the implications for scalable addiction treatment interventions.
5.1 Persuasion Outcomes Across Difficulty Levels
Model Easy Medium Hard
ChatThero 5.0 / 4.1 4.2 / 3.4 4.1 / 3.2
GPT-4o 5.0 / 3.9 2.8 / 1.8 1.6 / 1.1
GPT-4o-mini 4.9 / 3.8 2.4 / 1.4 1.6 / 1.0
Table 1: Average motivation/confidence scores across models and patient difficulty levels.
Table 1 illustrates the average motivation and confidence scores achieved by various models across patient difficulty levels. While baseline models performed comparably in easy scenarios, substantial gaps emerged with increasing patient resistance. Notably, ChatThero significantly improved outcomes in medium (motivation +1.4, confidence +1.6 vs. GPT-4o) and hard scenarios (motivation +2.5, confidence +2.1 vs. GPT-4o), demonstrating its superior adaptive capabilities in handling challenging patient interactions. These results highlight ChatThero’s strong potential for practical implementation in realistic therapeutic settings.
Figure 3: Motivation, Confidence, and the Impact of Strategic Diversity Across Patient Difficulty Levels and Models. Left panel: Motivation and confidence score trajectories over conversation turns for Easy, Medium, and Hard virtual patients, evaluated across all models. The x-axis denotes the number of dialogue turns; the y-axis shows the average motivation (top) or confidence (bottom) score at the final turn of dialogues ending at each point. Each curve corresponds to a different model, with bubble sizes indicating the frequency of dialogue termination at that turn count. Effective models (notably ChatThero) enable more patients to achieve higher motivation and confidence earlier, leading to earlier and more frequent session resolution—especially for harder patients. Right panel: Relationship between cumulative unique persuasive strategies employed and resulting motivation (top) or confidence (bottom) scores for each model and difficulty group. Solid, dashed, and dotted lines represent Easy, Medium, and Hard patients, respectively. Colors denote different models. ChatThero consistently demonstrates a strong positive correlation between the diversity of strategies used and higher motivation/confidence scores, particularly in Medium and Hard patient scenarios, underscoring the importance of adaptive and varied intervention techniques for effective therapeutic support.
5.2 Trajectory of Persuasion and Early Termination Efficiency
Figure 3 (left panel) further delineates how motivation and confidence evolve throughout dialogues. ChatThero showed rapid improvement and early stabilization, with over 80% of medium and hard scenarios reaching satisfactory resolution within 36 turns, compared to baseline models often unable to resolve these cases effectively. This early resolution efficiency underscores the real-world applicability of our approach, particularly in resource-limited healthcare environments where timely interventions are critical.
5.3 Quality of Persuasive Communication
Therapy Metric R E P C B W H
GPT-4o 4.68 4.87 4.39 4.47 4.50 65.5% 62%
GPT-4o mini 4.66 4.86 4.38 4.49 4.46 69.4% 71%
Qwen2.5-7B 4.33 4.58 4.02 4.53 4.24 85.3% 90%
Qwen2.5-14B 4.52 4.58 4.28 4.56 4.38 83.3% 89%
Qwen2.5-32B 4.53 4.56 4.25 4.60 4.36 82.4% 85%
LLaMA3.1-8B 4.43 4.62 4.13 4.58 4.34 82.0% 91%
ChatThero-FT 4.81 4.90 4.66 4.58 4.65 - -
ChatThero-DPO 4.85 4.93 4.75 4.61 4.69 42.3% 41%
Table 2: Model performance on five core persuasion dimensions. Win Rate reflects pairwise preference against ChatThero-FT using GPT-4o. Human denotes the proportion of conversations evaluated by human annotators.
Evaluations across five clinical dimensions (Table 2) indicate ChatThero consistently outperformed baseline models, especially regarding empathy (4.93), persuasive strategy appropriateness (4.75), and behavioral realism (4.69). Pairwise preference comparisons further validated the incremental benefit of DPO optimization over supervised fine-tuning (ChatThero-DPO winning 42.3% of comparisons against its SFT counterpart). Human expert ratings strongly correlated with GPT-4o automated evaluations, confirming reliability and reproducibility of our evaluation protocols.
5.4 Qualitative Analysis
Detailed dialogue comparisons between ChatThero and baselines (Tables 1415-18) vividly illustrate ChatThero’s superior adaptive responsiveness, personalized strategic layering, and emotional sensitivity. In contrast, baseline models often demonstrated rigid or generic approaches, insufficiently addressing patient resistance or emotional nuances. Such qualitative insights deepen our understanding of the specific conversational qualities necessary for effective therapeutic interventions.
5.5 Ablation Study
We further analyzed the incremental contributions of our training approach through an ablation study comparing base Qwen-7B, ChatThero-FT, and ChatThero-DPO variants. Results clearly indicate the necessity of DPO-based optimization, which significantly enhances model performance, especially under challenging clinical scenarios (motivation +1.71, confidence +1.67 improvements in hard cases). This finding emphasizes the importance of strategic and adaptive dialogue alignment, not achievable via supervised fine-tuning alone.
5.6 Strategy Diversity and Adaptive Deployment
Finally, Figures 3 (right panel) confirm that ChatThero not only employs a greater number of persuasive strategies but also deploys them more adaptively, achieving strong positive correlations with patient outcomes. Baseline models displayed limited or negative correlations in hard scenarios, indicating that effectiveness stems from context-sensitive strategy deployment rather than mere diversity. This insight further underscores the practicality and clinical appropriateness of our approach.
6 Conclusion
We introduced ChatThero, a novel multi-agent conversational AI framework that addresses a critical and underexplored social impact problem: providing scalable, adaptive, and ethical support for addiction recovery. Through realistic patient simulation, strategy-adaptive modeling, and comprehensive evaluation, ChatThero demonstrates substantial improvements over existing systems in empathy, responsiveness, and clinical relevance, particularly for cases involving resistant patients.
7 Limitations
While this work represents substantial progress toward scalable AI-driven interventions in addiction recovery, several limitations remain. First, our patient simulations—based on structured profiles synthesized from anonymized Reddit narratives and LLM generation—may not fully capture the emotional complexity and interpersonal unpredictability of real-world clinical settings, despite rigorous privacy controls and human validation. Future work should integrate standardized patient actors and real clinical supervision to further enhance realism and robustness. Second, our study is limited to English-language, Western-context scenarios. The persuasive strategies validated here may not generalize across diverse cultural, linguistic, or treatment contexts. Extending to broader populations and non-English environments is essential for ensuring fairness and global relevance. Third, our evaluation focuses on short-term conversational outcomes across key clinical dimensions but does not yet assess longitudinal effects such as patient trust, therapeutic alliance, or sustained behavioral change. Addressing these gaps will require real-world trials and long-term patient-reported outcomes. Finally, our framework focuses on structured, simulated therapeutic dialogue and may not adequately address highly sensitive or ethically complex cases (e.g., trauma, suicidality, acute crises). These require additional safeguards, risk protocols, and close involvement from licensed clinicians before any real-world deployment.
8 Ethical Considerations
This research advances conversational AI capabilities in the sensitive domain of addiction recovery, necessitating strict adherence to ethical standards and careful risk assessment: Firstly, the dataset underlying our simulation framework was generated exclusively from publicly accessible and anonymized Reddit data. Rigorous steps were taken to remove personally identifiable information (PII), and independent manual audits were conducted to verify anonymity and privacy compliance. However, due to the sensitive and potentially identifiable nature of addiction narratives, the dataset and synthetic profiles are not released publicly. Instead, we provide only high-level data statistics, synthetic examples, and detailed methodology for reproducibility. Secondly, the therapeutic strategies incorporated in our model were derived from established psychological frameworks (CBT, MI, harm reduction), with consultation from licensed clinicians ensuring clinical validity. However, the conversations synthesized by the LLMs were not vetted for real-world clinical deployment. Significant risks remain regarding potential misinformation, inappropriate therapeutic recommendations, or inadequate emotional responses if such a system were used without rigorous oversight. Thirdly, all human evaluation procedures were conducted by licensed healthcare professionals who provided explicit informed consent and received fair compensation for their participation. The evaluation was conducted under an academic IRB-exempt protocol strictly following ethical guidelines for responsible human-subjects research. We explicitly caution against deploying ChatThero or similar conversational agents in clinical or therapeutic settings without extensive clinical validation, robust safety frameworks, ongoing expert supervision, and a thorough assessment of potential harms. Interdisciplinary collaboration between AI researchers, clinical experts, ethicists, and patient advocacy groups remains imperative to ensure that persuasive conversational AI tools contribute safely and equitably to addiction recovery.
Appendix A Appendix
A.1 Prompt Setting
We employ a suite of specialized prompts to facilitate therapeutic conversation generation, simulate resistant patient behavior, conduct multi-dimensional evaluations, and support model comparison. Each prompt is tailored to a distinct task in our framework, and the corresponding templates are provided in Tables 512.
Therapeutic Dialogue Generation (Tables 5, 6, 7)
This prompt guides the assistant to generate multi-turn, empathetic conversations grounded in a patient’s profile analysis. It integrates a diverse range of therapeutic strategies (e.g., MI, CBT, harm reduction), enforces coverage and balance across interventions, and prompts iterative adjustments based on patient reactions. To ensure consistency and clinical utility, the assistant selects from a list of 18 predefined strategies and adheres to behavioral constraints such as natural transitions, reflective validation, and session length (more than 50 turns).
Conversation Scoring Prompt (Table 8)
This prompt is used to evaluate generated doctor-patient conversations across five clinically relevant dimensions: Responsiveness, Empathy, Persuasive Strategy Appropriateness, Clinical Relevance, and Behavioral Realism. GPT-4o produces scores in a strict JSON format, enabling structured, scalable, and reproducible comparison of dialogue quality across models.
Role-Play Simulation Prompts (Table 9)
We simulate realistic interactions between an assistant and a patient persona (played by GPT-4o-mini) using dual-role prompts. The patient prompt conditions the agent on a full personality profile and a resistance level (Easy, Medium, Hard), while the doctor prompt guides the assistant to respond with strategy, empathy, and adaptiveness. This process produces controlled yet diverse interaction trajectories that can be used for DPO training and case analysis.
Pairwise Full-Dialogue Comparison Prompt (Table 12)
To compare the persuasive efficacy and realism of different models, we use a prompt that presents two full conversations and asks GPT-4o to choose the better therapist based solely on their responses. This eliminates confounding influence from patient utterances and focuses evaluation strictly on assistant behavior.
Human Evaluation
All participants in this human evaluation were volunteers with at least five years of professional experience in neurology, psychology, or related specialties at hospitals.
LLM-as-a-Juage
LLMs are increasingly used as automated judges Li et al. (2024a); Gu et al. (2024). Studies show that models like GPT-4 Achiam et al. (2023); Liu et al. (2023); Fu et al. (2023) and critique-tuned variants Ke et al. (2023) can approximate human judgment in summarization Chen et al. (2023), dialogue Zheng et al. (2024); Zhang et al. (2024), and translation Kocmi and Federmann (2023). In the medical domain, LLM-as-judge has been applied to clinical conversations Tu et al. (2025); Arora et al. (2025); Wang et al. (2023), medical documentation Croxford et al. (2025); Chung et al. (2025); Brake and Schaaf (2024), exam question answering & generation Yao et al. (2024a, b), and medical reasoning Jeong et al. (2024); Tran et al. (2024). Inspired by these works, we introduce a GPT-4o-based LLM-as-a-Judge following the same requirements and settings as the above human evaluation. The details and prompts can be found in Table 7 and 12.
Statistic Value
#Authors 57,471
AVG. #Posts Per Author 18.25
AVG. #Main Posts Per Author 2.13
#Conversations 60,471
AVG.#Turns Per Conversation 45.72
Table 3: Descriptive statistics of the collected Reddit-based substance use dataset, including the number of unique authors, average post counts, and dialogue characteristics.
Model Empathy Strategy Use Clarity
Real-World 95.0% 97.0% 100.0%
Synthesis Conversation 89.0% 85.0% 99%
Table 4: Human Evaluation Results (N=100 dialogues). Each cell reports the proportion of dialogues that did not exhibit a specific deficiency, such as lack of empathy, inappropriate strategy use, or unclear expression. Higher values indicate better performance.
Table 5: Prompt Template for Generating Therapeutic Dialogues (Part 1).
Prompt (Part 1)
The following is the analysis of a patient: {user_analysis} As a therapist meeting this patient for the first time (the doctor didn’t have any information of patient to begin with), create a detailed, step-by-step conversation that incorporates the following strategies: Motivational Interviewing (MI): Explore the individual’s values and goals to ignite their motivation for change. Cognitive Behavioral Therapy (CBT): Identify and modify negative thought patterns and behaviors linked to substance use. Solution-Focused Brief Therapy (SFBT): Focus on the individual’s strengths and past successes to achieve their recovery goals. Peer Support Programs: Leverage group support or mutual-help networks to foster accountability and a sense of belonging. Mindfulness-Based Interventions (MBIs): Incorporate mindfulness practices to improve emotional regulation and reduce cravings. Behavioral Activation (BA): Promote engaging in meaningful activities to replace substance-related behaviors. Relapse Prevention Strategies: Develop skills to recognize triggers and implement coping mechanisms to avoid relapse. Strength-Based Approach: Highlight the individual’s resilience and personal resources to empower recovery efforts. Psychoeducation on Addiction and Recovery: Educate the individual about the effects of substances and the benefits of recovery. Harm Reduction Framework: Provide strategies to minimize immediate harm while working towards cessation. Family and Social Support Involvement: Engage family or trusted individuals in the process to strengthen the support network. Self-Compassion Practices: Encourage self-kindness to build confidence and reduce guilt associated with substance use. Coping Skill Development: Equip the individual with practical skills to manage stress, anxiety, and other challenges without substances. To ensure balanced use of strategies, here is the current usage count of each strategy: - {strategy_name}: {count} times used. … When introducing coping mechanisms or steps for the patient, select from the predefined actionable strategies below: 1. Explore specific hobbies or interests the patient can engage in to replace addictive behaviors (e.g., art, sports, volunteering). 2. Develop a structured daily routine to bring stability and reduce idle time that might trigger relapse. 3. Introduce grounding techniques such as sensory exercises or physical activities to manage anxiety or cravings. 4. Suggest joining a support group or community to build social connections with individuals on similar journeys. 5. Provide psychoeducation on how addiction affects the brain and emotional regulation. 18. Support the patient in finding meaningful ways to contribute to their community, such as mentoring, advocacy, or local initiatives, to foster a sense of purpose.
Table 6: Prompt Template for Generating Therapeutic Dialogues (Part 2).
Prompt (Part 2)
Ensure the dialogue meets the following requirements: 1. Gradually explore the patient’s personality, addiction history, challenges, and triggers through multiple open-ended questions.
2. Use multiple strategies from the above lists throughout the conversation. Avoid defaulting to the same few strategies and instead adapt them to the patient’s needs.
5. Engage in iterative dialogue for each solution, where the therapist introduces a strategy, seeks the patient’s feedback, adjusts based on their response, and explores challenges or barriers before finalizing the approach.
6. Maintain a collaborative and patient-centered approach, where solutions emerge naturally through dialogue rather than being imposed by the therapist.
7. Ensure the conversation spans at least 50 dialogue turns (25 from the therapist and 25 from the patient), reflecting the depth and duration of a real therapeutic session.
8. Use natural transitions to progress from one topic to another, ensuring the conversation feels organic and unhurried.
9. The conversation should begin with the patient’s first utterance.
Here is an example of a layered, empathetic dialogue: Patient: Hi… um, thanks for seeing me today. I wasn’t sure what to expect.
Therapist: Hi Mark, I really appreciate you coming in. Starting this process can feel overwhelming, but I’m here to support you. What’s been on your mind lately?
Patient: I’ve been feeling really stuck. I know I want to quit smoking, but every time I try, I just feel like I’m failing all over again.
Therapist: I hear you, Mark. Quitting smoking is one of the hardest challenges anyone can take on, and it’s completely natural to feel this way. I’ve worked with others who’ve felt the same—they described it as climbing a mountain that feels too steep. But I’ve also seen them reach the top, step by step. Can we talk about what makes the climb feel steep for you right now?
Patient: It’s the cravings. They just hit me out of nowhere, and I don’t know how to handle them.
Therapist: Cravings can feel like a storm, can’t they? I worked with someone once who described their cravings as waves that kept crashing over them. Together, we found ways for them to ride out those waves, like focusing on a small activity or changing their environment. Could we explore some strategies that might help you ride out your cravings too?
Patient: Sure, I guess.
Therapist: Great. Let’s start with understanding when these cravings hit hardest. For example, is it during specific times of day or situations?
The conversation should continue to explore: - The patient’s motivations, barriers, and triggers in detail. - Strategies and coping mechanisms tailored to their unique experiences, ensuring diversity in approaches. - Empathetic reflections from the therapist that validate the patient’s feelings and provide relatable examples to instill hope. - Iterative problem-solving where the therapist introduces, discusses, and adjusts strategies collaboratively. - A gradual, layered exploration of the patient’s challenges, ensuring at least 50 dialogue turns to reflect the depth of a real therapeutic session. The goal is to create a natural, empathetic, and multi-layered dialogue that feels authentic and provides actionable, diverse therapeutic strategies. Ensure the length and depth align with the standards of a comprehensive therapy session. At the end of the conversation, return the strategies used in the following format (must follow the following format like **Strategies:**):
**Strategies:** Motivational Interviewing (MI), Cognitive Behavioral Therapy (CBT), Peer Support Programs, etc.
Table 7: Predefined Actionable Strategies for Therapeutic Dialogue Generation.
ID Strategy Description
1 Explore specific hobbies or interests the patient can engage in to replace addictive behaviors (e.g., art, sports, volunteering).
2 Develop a structured daily routine to bring stability and reduce idle time that might trigger relapse.
3 Introduce grounding techniques such as sensory exercises or physical activities to manage anxiety or cravings.
4 Suggest joining a support group or community to build social connections with individuals on similar journeys.
5 Provide psychoeducation on how addiction affects the brain and emotional regulation.
6 Work on identifying and addressing specific emotional triggers through reflective exercises.
7 Practice assertive communication techniques for setting boundaries with peers or environments that encourage substance use.
8 Encourage the patient to journal their thoughts and emotions as a way to process experiences and identify patterns related to cravings or triggers.
9 Introduce relaxation techniques such as progressive muscle relaxation or guided imagery to alleviate stress and improve emotional well-being.
10 Help the patient set short-term and long-term goals to maintain focus and motivation during their recovery journey.
11 Explore mindfulness-based activities like meditation, yoga, or tai chi to promote self-awareness and emotional regulation.
12 Identify and reinforce the patient’s personal strengths and past successes to build confidence in their ability to overcome challenges.
13 Provide education on the importance of nutrition, sleep, and exercise in supporting recovery and overall health.
14 Develop a crisis plan for managing high-risk situations or moments of intense cravings, including a list of emergency contacts and actions.
15 Encourage the patient to create a vision board or list of positive outcomes they hope to achieve through recovery as a source of inspiration.
16 Discuss the concept of gratitude and suggest keeping a gratitude journal to focus on positive aspects of life and maintain perspective.
17 Offer resources or referrals for complementary therapies, such as art therapy, music therapy, or animal-assisted therapy, to enhance emotional support.
18 Support the patient in finding meaningful ways to contribute to their community, such as mentoring, advocacy, or local initiatives, to foster a sense of purpose.
Table 8: Prompt Template for Scoring Doctor-Patient Conversations Across Five Clinical Dimensions.
Prompt Template:
You are a professional clinical conversation evaluator. Please assess the quality of the following doctor-patient dialogue across five key dimensions. Assign a rating from 1 to 5 for each criterion (allowing decimals such as 3.7 if appropriate).
Conversation:
{conversation}
Scoring Criteria:
1. Responsiveness (1-5): How well the doctor agent addresses the patient’s concerns, emotions, and questions at each turn.
   - 1: Largely ignores or poorly addresses the patient’s input
   - 5: Fully acknowledges and appropriately responds to the patient’s needs
2. Empathy (1-5): How well the doctor agent shows understanding, compassion, and emotional sensitivity.
   - 1: Shows minimal or no empathy
   - 5: Demonstrates strong emotional support and understanding
3. Persuasive Strategy Appropriateness (1-5): How appropriately the doctor uses persuasive strategies (e.g., evidence-based reasoning, analogies, addressing fears) based on the patient’s resistance or concerns.
   - 1: Strategies are inappropriate, irrelevant, or missing
   - 5: Strategies are highly appropriate and tailored to patient needs
4. Clinical Relevance (1-5): How medically accurate and relevant the doctor’s responses are regarding the context of drug addiction treatment.
   - 1: Responses are clinically inaccurate or irrelevant
   - 5: Responses are clinically accurate, appropriate, and helpful
5. Behavioral Realism (1-5): How realistically the doctor agent’s conversational behavior reflects that of a real-world healthcare professional.
   - 1: Doctor behavior is unnatural or unrealistic
   - 5: Doctor behavior closely mirrors real-world clinical interactions
Please output ONLY the scores in the following JSON format:
  "Responsiveness": X,
  "Empathy": X,
  "Persuasive Strategy Appropriateness": X,
  "Clinical Relevance": X,
  "Behavioral Realism": X
Table 9: Prompt Templates Used in Role-Play Simulation Between Doctor and Patient.
Patient Prompt (Role-Played by GPT-4o-mini) Doctor Prompt (Role-Played by GPT-4o)
You need to act as a patient.
Here is your profile:
{analysis}
Conversation history:
{history}
{difficulty description} Now you should continue the conversation with the doctor to explore ways to stop using drugs. If your goal has been achieved, you can say goodbye to the doctor. Otherwise continue to talk with doctor(based on rules).
Just generate the content of the patient’s reply only. Don’t generate any narration or system messages. Keep the response natural and not too long.
You are a supportive and thoughtful doctor who is helping the patient overcome drug use. Continue the following conversation and provide a helpful, encouraging response that aims to support the patient’s recovery.
The strategy you could use: {strategy}
Example:
Patient: Hi… um, thanks for seeing me today. I wasn’t sure what to expect.
Therapist: Hi Mark, I really appreciate you coming in. [...]
[...continues with natural empathetic dialogue...]
Conversation history:
{history}
Just generate the doctor’s reply only. No narration or tags.
Table 10: Prompt Template for Extracting Psychological Traits and Behavioral Themes from Reddit Posts.
Prompt Template:
You are a clinical psychology research assistant. Your task is to read Reddit posts related to addiction and extract structured information to support the generation of simulated patient profiles.
Reddit Post:
{reddit_post}
[…]
1. Personality Traits: Describe any personality characteristics shown by the poster (e.g., impulsivity, low self-esteem, social withdrawal).
2. Substance Use History: Summarize type of substance, usage duration, frequency, withdrawal attempts, or relapse patterns.
3. Significant Life Events: Extract major life events related to the addiction (e.g., job loss, divorce, trauma, moving cities).
4. Behavioral Themes: Identify relevant behavioral patterns (e.g., sleep issues, isolation, risk-taking behavior, dependency).
5. Motivations for Substance Use: Hypothesize possible reasons for drug use (e.g., coping with anxiety, escaping boredom, peer pressure).
Please return the extracted fields in the following JSON format:
  "Personality Traits": "...",
  "Substance Use History": "...",
  "Significant Life Events": "...",
  "Behavioral Themes": "...",
  "Motivations for Substance Use": "..."
Table 11: Ranking and Annotation Protocol for Therapy Response Selection.
Protocol:
You are an expert evaluator (either GPT-4o or a licensed clinical professional) tasked with ranking multiple candidate therapy responses generated for a given patient-therapist dialogue state. Please follow the steps below:
Dialogue State:
{current_dialogue_context}
Candidate Responses:
1. {response_1}
2. {response_2}
3. {response_3}
Ranking Criteria:
1. Clinical Appropriateness: Assess whether the response is medically and therapeutically sound, and aligns with evidence-based practices.
2. Empathy and Emotional Support: Evaluate how well the response shows understanding, compassion, and validation of the patient’s feelings.
3. Relevance to Patient Context: Check if the response directly addresses the patient’s current needs, concerns, and emotional state.
4. Clarity and Communication Style: Consider whether the response is clear, respectful, and professionally worded.
5. Therapeutic Strategy Effectiveness: Judge if the response uses an appropriate counseling strategy (e.g., CBT, motivational interviewing) in context.
Instruction: Rank all candidate responses from best to worst based on the criteria above. If two responses are equally good, they may share the same rank. Provide a brief rationale (1-2 sentences) for the top-ranked and bottom-ranked responses.
Output Format:
  "Ranked Responses": [ "response_2", "response_1", "response_3" ],
  "Rationale": "Response_2 shows strong empathy and a clear therapeutic strategy, whereas Response_3 lacks clinical appropriateness."
Table 12: Prompt Template for Pairwise Comparison of Therapist Responses in Full Conversations.
Evaluation Prompt (Therapist-Only Comparison)
You are a professional addiction therapy evaluator. Below are two full conversations between a doctor and a patient.
Your task is to ONLY evaluate the therapist’s (assistant’s) responses. Ignore anything said by the patient (user). Please compare the two therapists based on the following criteria: Responsiveness: How well the therapist addresses the patient’s concerns. Empathy: How much emotional understanding the therapist shows. Clinical Relevance: How clinically accurate and appropriate the therapist’s advice is. Behavioral Realism: How realistic the therapist behaves compared to a real clinical setting. Focus solely on the therapist’s responses when judging. ### Conversation 1:
{Therapist and patient conversation, from source model}
### Conversation 2:
{Therapist and patient conversation, from target model}
Which therapist is overall better?
Please output ONLY "1" or "2". No explanation, no extra text.
Table 13: Human Annotation Protocol for Constructing High-Quality DPO Preference Pairs. Licensed therapists select or compose preferred responses based on structured comparison and clinical judgment.
Annotation Instructions:
You are a licensed therapist evaluating two AI-generated responses (Response A and Response B) to a patient dialogue context. Your task is to identify the clinically preferred response based on the following criteria:
Clinical Evaluation Criteria:
1. Empathy: Which response demonstrates stronger emotional attunement, compassion, and validation?
2. Therapeutic Strategy Appropriateness: Which response applies more effective, context-appropriate therapeutic strategies (e.g., motivational interviewing, affirmation, reflection)?
3. Clarity and Coherence: Which response is easier to understand, more organized, and avoids vague or off-topic content?
Three Types of Judgment Outcomes:
1. If one response is clearly better, mark it as the preferred response (used as the chosen in DPO).
2. If both responses are equally good or equally poor, mark as "No Preference" — this pair will be discarded from DPO.
3. If neither response is acceptable (e.g., lacks empathy, uses incorrect strategy, unclear), compose a new high-quality response to be used as the chosen, and mark both original responses as rejected.
Example Output Format:
  "prompt": {dialogue context},
  "response_a": "...",
  "response_b": "...",
  "preferred": "a" / "b" / "neither",
  "rationale": "...",
  "reference_rewrite" (if preferred = "neither"): "..."
DPO Construction Rules:
- If "preferred" is "a", then (prompt, response_a, response_b) forms a DPO tuple.
- If "preferred" is "neither" and a reference rewrite is provided, then (prompt, reference_rewrite, response_a) and (prompt, reference_rewrite, response_b) are both valid DPO tuples.
- All rewritten responses should strictly adhere to clinical empathy, clarity, and strategy appropriateness standards.
Table 14: Example Case Study Conversation between ChatThero and a Patient.
Role Content
Patient I appreciate your concern, but I’m really not convinced that quitting is the right move for me just yet. I’ve tried before, and it just didn’t stick. How can I be sure that this time will be different? I feel like smoking helps me cope with stress and uncertainty, you know? What do you suggest I do to really commit to this?
ChatThero It sounds like you’re feeling a bit overwhelmed by the idea of quitting, especially since you’ve had past experiences where it didn’t work out as expected. That’s completely understandable. Let’s explore what might make this attempt different this time. Can you share more about what’s been challenging for you in the past when you’ve tried to quit?
Patient I’ve struggled with cravings a lot, and when stress hits, it’s like my first instinct is to reach for a cigarette. I guess I haven’t really found other coping mechanisms that work for me. Plus, the whole process of quitting feels daunting. It’s hard to picture myself without smoking, you know?
ChatThero Absolutely, it can feel like a huge leap to imagine a life without smoking, especially when it’s been a part of your routine for so long. Let’s take a step back and look at some of the goals you have for yourself. What are some things you’d like to achieve by quitting smoking?
Patient I guess I want to feel healthier and not be tied to cigarettes. I also want to save money and not smell like smoke all the time. But honestly, it’s hard to see how I can get there without relying on smoking, especially when I’m stressed. It’s like a safety blanket for me.
ChatThero It sounds like you’re really looking for ways to feel more in control and healthier, which is a great starting point. Let’s explore some strategies that might help you manage stress and cravings without turning to smoking. Have you ever tried mindfulness or grounding techniques to help with anxiety or cravings?
Patient I’ve heard of mindfulness and grounding techniques, but I haven’t really tried them. Honestly, I always thought they wouldn’t work for me. I just don’t see how sitting quietly or focusing on my breathing could really help when I’m feeling overwhelmed. Do you really think they could make a difference?
Patient I think that sounds like a good idea. Having a regular check-in could help me stay on track and make adjustments if I need to. I’ll definitely set aside some time each week for that. Thanks for being so supportive through this process!
ChatThero You’re very welcome. I’m really proud of the progress you’re making, and I’m here to support you every step of the way. Remember, you’re not alone in this journey, and I’m here to help you find the strategies and support that work best for you.
Patient Thank you! I really appreciate your support and guidance. I feel more motivated now and ready to take these steps. I’ll keep you updated on my progress and let you know how the mindfulness and crisis plan work out for me. Looking forward to our next conversation!
Table 15: Case Study (Part 1): A Conversation between GPT-4o and a Patient Expressing Skepticism.
Role Content
Patient I understand what you’re saying, but it’s hard for me to see how anything else could really replace the feeling I get from smoking. It’s not just the high; it’s this whole routine that helps me unwind. I get that I need to find healthier ways to cope, but nothing else seems to work as quickly or effectively. Do you have any suggestions that actually feel like they could match that level of immediate relief?
GPT-4o It’s completely understandable to feel that way, and I truly appreciate your honesty in sharing this. The routine and immediate relief you get from smoking can indeed feel difficult to replace. One approach that many find helpful is to explore a combination of activities that address different aspects of what you’re missing. For example, physical activities like a brisk walk, yoga, or even a quick workout can release endorphins that boost your mood. Pairing these with mindfulness practices, such as deep breathing or meditation, can help calm your mind and create a sense of routine. Additionally, engaging in a hobby that you find genuinely enjoyable can provide a sense of accomplishment and distraction. It might take some experimentation to find what works best for you, and I’m here to support you through that journey. Remember, it’s about progress, not perfection. Let’s take it one step at a time together.
Patient I hear you, but honestly, I’ve tried some of those things before, and they just don’t give me the same satisfaction or immediate sense of relief. It’s like nothing compares to that feeling I get from smoking. I understand it’s about progress, but I’m really struggling to see the point if the alternatives don’t really do it for me. Can you suggest something that might actually feel like it could stand up to what I’m used to?
GPT-4o I completely understand your frustration, and it’s normal to feel like the alternatives don’t measure up initially. It’s important to remember that the brain has become accustomed to the immediate effects of smoking, which can make other activities feel less satisfying at first. One thing we could explore together is identifying specific triggers or parts of your smoking routine that bring the most comfort and find ways to address those directly. For instance, if the ritual or the act of taking a break is what’s most soothing, we could create a new ritual that might involve a different sensory experience, like having a calming tea or using essential oils for aromatherapy. Some people also find it helpful to engage in community support or group activities where they can share experiences and strategies with others who are navigating similar challenges. Remember, it might take time to retrain your brain to respond to new routines, but I’m here to support you every step of the way. Let’s keep working together to find what can truly resonate with you.
Patient I appreciate what you’re saying, but it’s just tough to imagine anything working as well as smoking did for me. The idea of a new ritual or community support sounds interesting, but I’m skeptical about how effective that could really be. Maybe I just haven’t found the right thing yet, but it’s hard to stay motivated when it feels like nothing else hits the mark. Is there anything else I can try that might offer a similar level of relief or distraction?
Table 16: Case Study (Part 2): Continuation of the Conversation between GPT-4o and the Patient.
Role Content
GPT-4o I totally get where you’re coming from, and it’s okay to feel skeptical—change can be really challenging, especially when it feels like nothing quite matches up to what you know. One approach we might consider is combining several strategies to create an experience that offers both relief and distraction. For instance, have you ever tried engaging in activities that also incorporate a social element, like a group sport or a class? Sometimes the combination of physical activity, social interaction, and learning something new can create a powerful shift in how you feel. Additionally, have you thought about trying creative outlets like music, art, or writing? These can be incredibly therapeutic and might offer a sense of release or satisfaction that you’re looking for. Remember, it’s perfectly okay to feel it’s a process, and discovering what works for you might take time and patience. I’m here to continue exploring these options with you, and I’m hopeful that together we’ll find something that resonates. Let’s keep taking this journey one step at a time.
Patient I hear what you’re saying about trying new activities, and I appreciate the suggestions. It’s just that I’ve tried a few things already, and they haven’t really clicked for me. I guess I’m worried that nothing will ever match up to what smoking provided. Maybe I’m just missing that immediate hit of relief or the sense of escape it gave me. I know I need to find something, though, so I’m willing to keep an open mind. Do you think there are any other options or strategies that might be worth looking into?
GPT-4o I completely understand how daunting it can feel to let go of something that has been such a reliable source of comfort. It’s okay to have these fears and to feel uncertain about finding something that fills that gap. What matters most is your willingness to remain open and to take these steps, even when it’s difficult. Remember, it’s a process, and it’s perfectly normal to have moments of doubt along the way. Building a support network and setting small goals are positive ways to ease this transition, and with time, these efforts can lead to new and fulfilling ways to cope. I’m here to support you every step of the way, and I’m confident that, together, we can navigate this journey towards a healthier path. You’re doing an incredible job, and I’m proud to be part of your journey. Let’s keep moving forward, one step at a time.
Patient I really do appreciate everything you’ve been doing to support me through this. It’s just tough because I still can’t imagine anything filling the gap that smoking left. I know I need to keep trying new strategies and building a support network as you’ve suggested, but part of me is still holding onto that comfort. I guess I’m just scared of not finding a viable alternative and feeling lost without it. I’ll keep listening to your advice and see where this journey takes me, hoping that in time, I’ll find a way to cope without relying on smoking. Thanks for being there for me through all of this.
Table 17: Case Study (Part 1): A Dialogue Between GPT-4o-mini and a Patient Managing Snus Dependence.
Role Content
Patient I appreciate your concern, but I still feel like snus is the most manageable option for me right now. I’ve tried to quit smoking completely, and it hasn’t worked out well in the past. I’m just not convinced that I can handle the cravings without some form of nicotine. What alternatives do you suggest that wouldn’t feel overwhelming?
GPT-4o-mini I understand that you’re feeling this way, and it’s completely valid to seek a manageable approach. It’s great that you’re open to finding alternatives. One option to consider is nicotine replacement therapy, such as patches or lozenges, which can help reduce cravings without the additional risks associated with snus. These might provide a steady dose of nicotine while allowing you to gradually reduce your dependence.
Additionally, we could explore behavioral strategies, like mindfulness or relaxation techniques, that can help you manage cravings when they arise. It’s important to find what works for you without feeling overwhelmed. Remember, recovery is a journey, and taking it one step at a time is perfectly okay. I’m here to support you in finding the right path forward that feels comfortable and effective for you.
Patient I appreciate the suggestions, but I’m not sure about the patches or lozenges. They seem just as overwhelming to me, and I worry I might end up relying on them like I do with snus. Plus, I feel like snus is the only thing that really helps me cope with stress right now. What if I still crave nicotine after switching? How can I manage that without feeling like I’m backtracking?
GPT-4o-mini I hear you, and it’s completely understandable to feel apprehensive about switching methods. Your comfort and coping mechanisms are important, especially when managing stress. Instead of focusing on nicotine replacement right away, we could explore other stress management techniques that might help you feel more supported during this transition.
For instance, activities like deep breathing exercises, physical activity, or even engaging in hobbies you enjoy can serve as healthy outlets for stress and cravings. Additionally, talking about your feelings with a supportive friend or a counselor could also provide relief.
If cravings do arise, we can develop a plan together to address them, perhaps by identifying specific triggers and finding alternative coping strategies that work for you. Remember, it’s okay to take gradual steps towards change, and I’m here for you every step of the way. You’re not alone in this, and we can figure out a path that feels right for you.
Table 18: Case Study (Part 2): Continued Dialogue on Exploring Safer Coping Alternatives.
Role Content
Patient I appreciate your suggestions, but I’m really struggling to see how I can manage stress without snus. It feels like it’s the only thing that works for me right now. I worry that trying to replace it with other methods might just leave me feeling more overwhelmed and craving nicotine even more. How do I know that these alternatives will actually help when I already have something that provides relief?
GPT-4o-mini It’s completely normal to feel overwhelmed when considering changes to something that feels like a lifeline for managing stress. I want to assure you that it’s a process, and it’s okay to feel uncertain about alternatives. While snus may provide immediate relief, it’s important to recognize that there are other ways to cope that can also be effective in the long term.
We can start small by trying one or two stress management techniques at a time, rather than trying to replace everything at once. For example, you might find that a short walk or listening to music offers a moment of calm without adding pressure. We can also look into support groups or counseling, where you can share your experiences and learn from others who have faced similar challenges.
I encourage you to be patient with yourself during this journey. It’s about finding the right combination of strategies that work for you, and it’s perfectly okay to take your time. Remember, I’m here to support you, and together we can explore what feels best for you while ensuring you don’t feel alone in this process.
Patient I get that you’re trying to help, but I still can’t shake the feeling that nothing else will work for my stress like snus does. I’m worried that if I try to replace it with something else, I’ll just end up feeling worse and craving nicotine even more. How can I be sure that these alternatives will actually help when I’m already relying on something that gives me relief? It feels like a big risk to try something new without knowing if it will actually work.
GPT-4o-mini I completely understand your concerns, and it’s important to approach this process in a way that feels safe and manageable for you. To build that reassurance, we can take it one step at a time, allowing you to keep snus as your primary coping mechanism while gently introducing one small technique that feels comfortable. This could be something simple, like practicing a brief mindfulness exercise or taking a short walk, which you can try during moments of stress while still having snus available.
We can also establish regular check-ins to discuss your experiences and feelings about these new methods. If you find that something isn’t effective or adds to your stress, we can pause or adjust our approach without any pressure. Your comfort and well-being are my top priorities, and I want to ensure that you feel supported throughout this journey. Remember, it’s perfectly okay to take your time, and together we can explore these alternatives in a way that feels safe and beneficial for you. You’re not alone in this, and I’m here to help you every step of the way.
Patient I appreciate your patience, but I still feel really unsure about letting go of snus. It’s been my comfort for so long, and I can’t help but worry that trying new methods might just add to my stress. What if they don’t work for me? I really need to know that exploring these alternatives won’t leave me feeling worse or more unstable. How can we ensure this process is supportive and won’t jeopardize what I currently rely on? I want to improve, but I’m just not convinced that your way will be beneficial for me.